# Methods and Trends in Detecting Generated Images: A Comprehensive Review

### v1

@@ 3.0.7. GenImage by Zhu: 2024
Zhu et al. (Zhu et al., 2024) introduced a large-scale dataset containing synthetic images generated by advanced state-of-the-art GANs and diffusion models (Wukong (Wukong, 2022) and VQDM (Gu et al., 2022), including commercial models, alongside real images. The dataset comprises approximately 1.3 million synthetic images and 1.3 million real images, covering diverse general image content. The authors also propose two real-world analysis factors to assess detection performance: (a) Cross-Generator Image Classification: training detectors on images generated by one model and evaluating them on images from different generators, and (b) Degraded Image Classification: testing detectors on degraded images affected by factors such as low resolution, JPEG compression, and Gaussian blur.
3.0.8. CIFAKE Dataset by Bird and Lotfi: 2024
@@
Bird and Lotfi (Bird and Lotfi, 2024) released a dataset comprising 120,000 images, equally divided between 60,000 real images and 60,000 synthetic images. The real images originate from CIFAR-10 (Krizhevsky et al., 2009), which includes ten object categories: airplane, automobile, bird, cat, deer, dog, frog, horse, ship, and truck. The synthetic images were generated using diffusion models, including commercially available ones.
3.0.9. AIGCDetection Benchmark Dataset by Zhong et al.: 2024
@@
Zhong et al. (Zhong et al., 2024) introduced a curated benchmark dataset incorporating images generated by 17 different generative models, including GANs and diffusion models. The dataset aggregates samples from ForenSynths (Wang et al., 2020b) and GenImage (Zhu et al., 2024), and includes additional images from recent diffusion models, offering a comprehensive benchmark for generative image forensics.
3.0.10. ImagiNet Dataset by Boychev: 2024
@@
Boychev et al. (Boychev and Cholakov, 2024) presented the ImagiNet dataset, comprising 200,000 images equally split between real and synthetic categories. The dataset further categorizes images into photos, paintings, faces, and uncategorized types. The synthetic images are generated using GANs, diffusion models, and proprietary generative models. For evaluation, the dataset is divided into 160,000 training images and 40,000 test images.
3.0.11. Chameleon Dataset by Yan et al.: 2024
@@
Yan et al. (Yan et al., 2024) introduced the Chameleon dataset, which focuses on realistic AI-generated image detection. The dataset consists of 150,000 synthetic images covering content categories such as humans, animals, scenes, and objects, generated using GANs and diffusion models. In addition, it includes 20,000 real images to facilitate forensic evaluation.
3.1. Discussion
@@
(10) $\text{Acc}=\frac{TP+TN}{TP+TN+FP+FN}$
(11) $\text{AP}=\sum_{n}(R_{n}-R_{n-1})P_{n}$
where $TP$, $TN$, $FP$, and $FN$ represent true positives, true negatives, false positives, and false negatives, respectively. The AP is computed as the weighted sum of precision $P_{n}$ at different recall levels $R_{n}$.
For clarity, we present a comparative analysis of methods across five distinct categories, following a general review of their reported performances in the literature. The results are detailed in Table 3, Table 6, and Table 8. Additionally, a comprehensive performance comparison across all five categories using the UnivFD dataset (Ojha et al., 2023) is provided in Table 9 for classification accuracy and Table 10 for average precision. The evaluation results on the UnivFD dataset (Ojha et al., 2023) are selected due to its inclusion of state-of-the-art detection methods, particularly those utilizing Vision-Language approaches, which have demonstrated strong forensic capabilities in synthetic image detection. These methods exhibit significant potential for real-world applications, reinforcing their role in advancing forensic detection.
4. Conclusion
The advancement of generative AI in image synthesis and forensic detection has been an ongoing and increasingly active research area, driven by both the successes of AI and the ethical concerns it introduces. From a practical standpoint, forensic detection must remain ahead to mitigate potential human-related harm. While several reviews exist in the current literature, they do not comprehensively cover detection methods capable of identifying images generated by state-of-the-art generative models, nor do they adequately address the role of multimodal approaches in detection. To bridge this gap, we have conducted a comprehensive survey of detection methodologies. First, we categorize these methods into five primary groups based on their underlying approaches. We then analyze their comparative performance on publicly available datasets and assess whether they satisfy three key criteria for evaluating their generalizability. Our findings indicate that detection methods leveraging multimodal frameworks tend to exhibit greater robustness and adaptability across different generative models. While this study primarily focuses on the image domain, the insights gained, specifically regarding frequency-domain analysis, can be extended to forensic detection in other modalities, such as audio and video analysis. Transforming such data into the frequency domain with appropriate operations may reveal anomalies that serve as modality-specific fingerprints for adversarial detection.
Acknowledgment
This material is based in part upon work supported by the National Science Foundation under Grant Nos. MRI20 CNS-2018611 and MRI CNS-1920182.

### v2

@@ 4.0.7 GenImage by Zhu: 2024
Zhu et al. Zhu et al. (2024) introduced a large-scale dataset containing synthetic images generated by advanced state-of-the-art GANs and diffusion models (Wukong Wukong (2022) and VQDM Gu et al. (2022), including commercial models, alongside real images. The dataset comprises approximately 1.3 million synthetic images and 1.3 million real images, covering diverse general image content. The authors also propose two real-world analysis factors to assess detection performance: (a) Cross-Generator Image Classification: training detectors on images generated by one model and evaluating them on images from different generators, and (b) Degraded Image Classification: testing detectors on degraded images affected by factors such as low resolution, JPEG compression, and Gaussian blur.
4.0.8 CIFAKE Dataset by Bird and Lotfi: 2024
@@
Bird and Lotfi Bird and Lotfi (2024) released a dataset comprising 120,000 images, equally divided between 60,000 real images and 60,000 synthetic images. The real images originate from CIFAR-10 Krizhevsky et al. (2009), which includes ten object categories: airplane, automobile, bird, cat, deer, dog, frog, horse, ship, and truck. The synthetic images were generated using diffusion models, including commercially available ones.
4.0.9 AIGCDetection Benchmark Dataset by Zhong et al.: 2024
@@
Zhong et al. Zhong et al. (2024) introduced a curated benchmark dataset incorporating images generated by 17 different generative models, including GANs and diffusion models. The dataset aggregates samples from ForenSynths Wang et al. (2020b) and GenImage Zhu et al. (2024), and includes additional images from recent diffusion models, offering a comprehensive benchmark for generative image forensics.
4.0.10 ImagiNet Dataset by Boychev: 2024
@@
Boychev et al. Boychev and Cholakov (2024) presented the ImagiNet dataset, comprising 200,000 images equally split between real and synthetic categories. The dataset further categorizes images into photos, paintings, faces, and uncategorized types. The synthetic images are generated using GANs, diffusion models, and proprietary generative models. For evaluation, the dataset is divided into 160,000 training images and 40,000 test images.
4.0.11 Chameleon Dataset by Yan et al.: 2024
@@
Yan et al. Yan et al. (2024) introduced the Chameleon dataset, which focuses on realistic AI-generated image detection. The dataset consists of 150,000 synthetic images covering content categories such as humans, animals, scenes, and objects, generated using GANs and diffusion models. In addition, it includes 20,000 real images to facilitate forensic evaluation.
4.1 Metrics
@@
$\text{Acc}=\frac{TP+TN}{TP+TN+FP+FN}$ (6)
$\text{AP}=\sum_{n}(R_{n}-R_{n-1})P_{n}$ (7)
where $TP$, TNTN, FPFP, and $FN$ represent true positives, true negatives, false positives, and false negatives, respectively. The AP is computed as the weighted sum of precision $P_{n}$ at different recall levels $R_{n}$.
5 Conclusion
The advancement of generative AI in image synthesis and forensic detection has been an ongoing and increasingly active research area, driven by both the successes of AI and the ethical concerns it introduces. From a practical standpoint, forensic detection must remain ahead to mitigate potential human-related harm. While several reviews exist in the current literature, they do not comprehensively cover detection methods capable of identifying images generated by state-of-the-art generative models, nor do they adequately address the role of multimodal approaches in detection. To bridge this gap, we have conducted a comprehensive survey of detection methodologies. First, we categorize these methods into seven primary groups based on their underlying approaches. We then describe each method in detail, followed by an analysis of their comparative performance on publicly available datasets, and assess whether they satisfy three key criteria for evaluating their generalizability. Our findings indicate that detection methods leveraging multimodal frameworks tend to exhibit greater robustness and adaptability across different generative models. Building upon this comparative analysis, we observe that spatial- and frequency-domain methods form the foundational layer of forensic understanding by identifying low-level pixel and spectral inconsistencies, yet they often struggle under compression or strong generative realism. Fingerprint- and patch-based methods address these weaknesses through architecture-agnostic learning and localized texture analysis, improving sensitivity and cross-generator robustness. Training-free approaches further enhance scalability by eliminating retraining needs, although interpretability and adversarial resilience remain open challenges. Finally, multimodal and reasoning-based models, particularly those integrating vision-language and large language models, bring explainability and semantic grounding but at the expense of heavy computational requirements. Looking ahead, the convergence of these paradigms offers a promising pathway. Hybrid frameworks that combine the reasoning and semantic grounding of multimodal models with the lightweight efficiency of training-free detection could lead to robust, interpretable, and real-time forensic systems. Future research should also emphasize cross-modal benchmarking and unified evaluation standards. Through these directions, the field can advance toward trustworthy and transparent generative AI systems capable of safeguarding authenticity in the digital era.
[…]